# Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models

### v1

@@ Abstract
Large language models (LLMs) exhibit remarkable capabilities in natural language processing but face catastrophic forgetting when learning new tasks, where adaptation to a new domain leads to a substantial decline in performance on previous tasks. In this paper, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. Aiming to reduce the scale of output change while introduce minimal constraint on model capacity, CLoRA imposes constraint on the direction of updating matrix’s null space. Experimental results on commonly used LLM finetuning tasks reveal that CLoRA significantly outperforms existing LoRA subsequent methods on both in-domain and out-domain evaluations, highlighting the superority of CLoRA as a effective parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
@@ 1 Introduction
The dimension of the null space for the updating matrix is directly determined by the rank of it, LoRA already addressed. A key factor remains, the direction of null space, which influence input components that fall in, but free-learned LoRA does not constraint. CLoRA constraint the direction of null space of updating matrix by introducing a pre-defined subset of that, this is implemented by orthogonal regularization with a pre-defined matrix. Unlike methods that impose restrictions on rank or norm, which significantly influence model capacity, CLoRA introduces constraint on the direction of the null space. Experimental results on commonly used LLM finetuning evaluations demonstrate that our proposed CLoRA outperforms existing methods on both in-domain downstream tasks and out-domain evaluations. This indicates the superiority of CLoRA as an effective approach for parameter-efficient finetuning with catastrophic forgetting mitigating. Additionally, we take analysis on parameters of the learned model, results show that CLoRA reduces the scale of output change with minimal impact on model capacity.
@@
  • We propose CLoRA, a subspace regularization method on LoRA structure, which serves as an advanced parameter-efficient finetuning technique with catastrophic forgetting mitigating for LLMs.
  • Our proposed CLoRA demonstrates superior performance on commonly used LLM finetuning tasks compared to existing methods. It has also shown remarkable mitigating of catastrophic forgetting in out-domain evaluations.
@@ 2 Related Works
2.1 Reducing Catastrophic Forgeting
Catastrophic forgetting is a significant challenge in various transfer learning scenarios, including continual learning Wang et al. (2023a) and LLM continued training Wu et al. (2024). In these settings, continued training on new tasks may impair abilities of the pre-trained model. Approaches for mitigating catastrophic forgetting can be broadly categorized into data-based, architecture-based and learning-based methods.
@@ 2.1 Reducing Catastrophic Forgeting
performs continued training in a shared vector space, controlling learning process by adding regularization term on loss or applying specific optimization designs. Orthogonal Gradient Descent (OGD) Farajtabar et al. (2020) constrains the parameters to move within the orthogonal space defined by the gradients of previous tasks. Continual LoRA (C-LoRA)Smith et al. (2024) and O-LoRA Wang et al. (2023b) introduce regularization with previous continual learned parameters for reducing interference in the multi-stage training setting. Our proposed CLoRA imposes orthogonal regularization similar to O-LoRA, but the regularization matrix is not restricted to be the previous learned parameter, thus CLoRA can be used for one-stage continued training whereas O-LoRA cannot.
@@ 3.2 Problem Definition
We aim to mitigate catastrophic forgetting in LLM continued training. Catastrophic forgetting menifest as performance decline on tasks from previous domain when training on new domain. To evaluate this, we conduct both in-domain tasks (demonstrating the effectiveness of training) and out-domain tasks (from previous domain, indicating the degree of forgetting). We conduct experiments in the one-stage tuning setting, specifically, we finetune a base LLM on one training dataset, then take in-domain and out-domain evaluations. Note that there is no clear domain specific for base LLMs, but benchmarks exist for evaluating the ability of LLMs on wide range of domains Gao et al. (2024), and we take those with minimal overlap with training data for out-domain evaluation.
@@ 4 Method
For the CLoRA regularization matrices, we suggest that they should have othogomal rows. Specifically, for regularization matrix $P\in\mathbb{R}^{m\times k}$, $||P[:,i]||=1$ for every $i$, and $P[:,i]P[:,j]=0$ for $i\neq j$. This ensures uniform regularization. We propose several variants as follows,
  • Random Initialize: Initialize with an orthogonal matrixSaxe et al. (2014).
  • Initialize from SVD: Previous works Sharma et al. (2024); Wang et al. (2024a) have explored the roles of singular value decomposition (SVD) components in LLM parameters. We adopt this intuition to initialize CLoRA regularization matrices from SVD. For a SVD decomposition of parameter $W=USV^{T}$ with rank $r$, where $W\in\mathbb{R}^{m\times n}$, $U\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix, $V\in\mathbb{R}^{n\times r}$. For CLoRA updating $\Delta W=AB^{T}$, we initialize the regularization matrix $P_{A}\in\mathbb{R}^{m\times k}$ as $U[:,\mathbf{s}]$, $P_{B}\in\mathbb{R}^{n\times k}$ as $V[:,\mathbf{s}]$, where $\mathbf{s}$ is a list of selecting index with length $k$.
5 Experiment and Analysis
In this section, we conduct experiments to evaluate and analyze our proposed CLoRA method. All evaluations are available in lm-eval Gao et al. (2024) with MIT License. We aim to answer these research questions,
  • RQ1: Does CLoRA perform effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating?
  • RQ2: How do the size and selection of regularization matrices influence the performance of CLoRA?
@@ 5 Experiment and Analysis
5.1 Evaluating for Tuning on Commonsense Reasoning Dataset (RQ1)
To answer RQ1, we conduct experiments following the setting in the previous works on PEFTLiu et al. (2024); Wang et al. (2024a). Specifically, we finetune LLaMA-2-7B Touvron et al. (2023) with a dataset for commonsense reasoning. We evaluated the model on both in-domain downstream tasks to evaluate learning effectiveness and on out-domain tasks with minimal overlap with the training dataset to measure the degree of forgetting. We report partial results here, full results on each CLoRA variants and results on finetuning LLaMA-3-8BAI@Meta (2024) are reported in Appendix B.
5.1.1 Experiment Setup
Datasets and Tasks
We use Commonsense170K Hu et al. (2023) for finetuning. For in-domain evaluation, eight commonsense reasoning datasets are used, including BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag(HS) Zellers et al. (2019), WinoGrande(WG) Sakaguchi et al. (2019), ARC-e, ARC-c Clark et al. (2018), and OBQA Mihaylov et al. (2018). The tasks are formulated as multiple-choice problem, and we report accuracy based on the last checkpoint. For out-domain evaluations, BIG-Bench-Hard(BBH) Suzgun et al. (2022) and MMLU-Pro Wang et al. (2024b) are used. These benchmarks encompass challenging subsets of tasks across a wide range of domains and are widely employed for evaluating the capabilities of LLMs. Additionally, they include samples that are more complex than those in our training data, ensuring minimal overlap. We use lm-eval Gao et al. (2024) for reporting out-domain evaluation.
Comparison Methods
@@ 5.1.1 Experiment Setup
  • CLoRA: Random initialize variant of CLoRA, it is the basic setting of our method. We will include results for other variants in the subsequent section 5.2.
Category Method BoolQ PIQA SIQA HS WG ARC-e ARC-c OBQA Avg.
ChatGPT ChatGPT 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9
LoRAs DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0
Reducing LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8
Forgetting LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8
LoRA-L2 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8
Ours CLoRA 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7
@@
Table 2: In-domain results on commonsense reasoning evaluations, with accuracy scores (%) reported. Bold font indicates the highest performance for each dataset across the different PEFT methods.
Hyperparameter
We use the same hyperparameter configurations as Wang et al. (2024a). Details are listed in Appendix A. Notably, we use 32 for updating matrix rank $r$ as default for all methods. For CLoRA regularization matrix, we select $k$ in $[128,256,512,1024,2048]$. For LoRA-L2, 1e-5 is used for weighting of L2 regularization, we note that 1e-4 is also tested, but too large for getting effective tuning.
5.1.2 Results and Analysis
In-domain Evaluation
We report in-domain results in Table 2. LLaMA-2-7B cannot be directly applied to these tasks without finetuning, following previous worksWang et al. (2024a), we include the ChatGPT baseline reported in Liu et al. (2024) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought Wei et al. (2022). For CLoRA, we report the best result across all results for all $k$.
For in-domain evaluation, results demonstrate that CLoRA outperforms on all datasets, surpassing the best baseline by 2.9 points in average accuracy (%). These outcomes suggest that, although proposed primarily for mitigating catastrophic forgetting, CLoRA also performs as a effective PEFT method. We attribute this to the nature of LLM finetuning, as an instance of transfer learning, the performance of LLM finetuning is strongly correlated with the base model’s ability, when catastrophic forgetting occurs during training, the base model’s strength may diminish. Therefore, we hypothesize that a method with effective capacity-forgetting balancing would exhibit strong effectiveness in LLM finetuning.
Method BBH MMLU-Pro
base 34.91 18.56
PiSSA 29.54 11.33
DoRA 28.24 11.67
LoRA 26.69 14.46
MiLoRA 25.14 17.74
LoRA-r8 26.90 14.58
LoRA-r16 26.73 11.54
LoRA-L2 32.93 16.59
CLoRA 38.67 20.59
Table 3: Out-domain results on LLM benchmarks, with accuracy scores (%) reported.
12825651210242048767676767878787880808080828282828484848486868686kAvg. Accuracy (%)CLoRA-randomCLoRA-majorCLoRA-minor
1282565121024204820202020222222222424242426262626282828283030303032323232kAvg. Accuracy (%)CLoRA-randomCLoRA-majorCLoRA-minor
Figure 3: Performace for CLoRA on commonsence reasoning finetuning setting in-domain(above), out-domain(below) with different variant and k.
Out-domain Evaluation
We report out-domain results for CLoRA and baselines in Table 3. “base” denotes the result on the base model LLaMA-2-7B, “CLoRA” denotes the result of best model from in-domain evaluation.
In out-domain evaluation, all baselines underperform compared to the base model, highlighting the severe issue of catastrophic forgetting in this experimental setup. Notably, our proposed CLoRA not only outperforms all baselines, but also surpasses the base model’s performance. We attribute this to the effective capacity-forgetting balancing for CLoRA, enabling the effective extraction of generally useful knowledge from the commonsense reasoning training dataset.
Answer for RQ1:
The superior performance in both in-domain and out-domain evaluations demonstrates that our proposed CLoRA serves effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating.
5.2 Evaluating for different CLoRA variants and hyperparameters (RQ2)
To answer RQ2, we compare results for various CLoRA variants to demonstrate the impact of regularization matrix’s size and selection.
Datasets and Tasks
Same as in section 5.1, we finetune LLaMA-2-7B on Commonsense170K, and conduct in-domain evaluation on eight commonsense reasoning tasks and out-domain evaluation on two LLM benchmarks. We report average results for each evaluation. To ascertain whether the optimal $k$ value is consistent across tasks, we additionally perform an experiment in the math domain. we finetune LLaMA-2-7B on the MetaMathQA Yu et al. (2024), which contains 395K samples augmented from the training set of GSM8K Cobbe et al. (2021) and MATHHendrycks et al. (2021). We use test set of GSM8K and MATH for evaluation and report the results on the last checkpoint.
Comparision Methods
  • Different $k$(-k*): We test CLoRA with different regularization matrix size ($k$). Higher $k$ imposes stricter constraints on the updating matrix.
  • CLoRA-random: CLoRA with random initialization.
For math evaluation, PiSSAMeng et al. (2024) and MiLoRAWang et al. (2024a) are included as baselines, we use CLoRA-major for CLoRA variant.
5.2.1 Results and Analysis
Commonsense Reasoning Setting
We illustrate the performance for in-domain and out-domain evaluations with different $k$ for commonsense reasoning setting as in Figure 3. The results indicate that for all CLoRA variants tested, as $k$ increases, the scores of both in-domain and out-domain evaluation generally exhibit an upward trend. Despite similar performance across variants under the same $k$, CLoRA-major generally performs better, highlighting the advantage of preserving the well-learned major components Wang et al. (2024a) of the base model.
Math Setting
Results on math setting are reported in Table 4. CLoRA outperforms other baselines, and the optimal $k$ value found to be around 128, differing from the 2048 in commonsense reasoning setting. This discrepancy can be attributed to the complexity of math tasks, which necessitate greater model capacity during finetuning. These findings suggest that adjusting $k$ benefits the balancing of capacity and forgetting, demonstrating the flexibility of our approach.
@@ 5.2.1 Results and Analysis
Answer for RQ2:
Results indicate that the choice of regularization matrix does influence the effectiveness of CLoRA, albeit not significantly. Generally, we recommend using random initialization (CLoRA-random) or initialization from major SVD components (CLoRA-major). Regarding the choice of $k$, as k increases, a trend of performance improvement followed by a decline is observed. Different tasks may correspond to different optimal $k$, we suggest that a smaller $k$ could be used for more challenging tasks.
5.3 Understanding Capacity-Forgetting Balancing(RQ3)

### v2

@@ Abstract
Large language models (LLMs) exhibit remarkable capabilities in natural language processing but face catastrophic forgetting when learning new tasks, where adaptation to a new domain leads to a substantial decline in performance on previous tasks. In this paper, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. Aiming to reduce the scale of output change while introduce minimal constraint on model capacity, CLoRA imposes constraint on the direction of updating matrix’s null space. Experimental results on one-stage LLM finetuning tasks and continual learning settings highlight the superority of CLoRA as a effective parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
@@ 1 Introduction
The dimension of the null space for the updating matrix is directly determined by the rank of it, which LoRA already addressed. A key factor remains, the direction of null space, which influence input components that fall in, but free-learned LoRA does not constraint. CLoRA constraint the direction of null space of updating matrix by introducing a pre-defined subspace, this is implemented by orthogonal regularization with a pre-defined matrix. Unlike methods that impose restrictions on rank or norm, which significantly influence model capacity, CLoRA introduces constraint on the direction of the null space. We take experiments on commonly used one-stage LLM finetuning evaluations and continual learning evaluations, results indicate the superiority of CLoRA as an effective approach for parameter-efficient finetuning with catastrophic forgetting mitigating. Additionally, we take analysis on parameters of the learned model, results show that CLoRA reduces the scale of output change with minimal impact on model capacity.
@@
  • We propose CLoRA, a subspace regularization method on LoRA, which serves as an advanced parameter-efficient finetuning technique with catastrophic forgetting mitigating for LLMs.
  • Our proposed CLoRA demonstrates superior performance on both in-domain and out-domain evaluation in commonly used one-stage LLM finetuning setting. Additionally, it showns remarkable mitigating of catastrophic forgetting in continual learning setting.
@@ 2 Related Works
2.1 Mitigating Catastrophic Forgeting
Catastrophic forgetting is a significant challenge in various transfer learning scenarios, including continual learning Wang et al. (2023a) and LLM finetuning Wu et al. (2024). In these settings, continued training on new tasks may impair abilities of the pre-trained model. Approaches for mitigating catastrophic forgetting can be broadly categorized into data-based, architecture-based and learning-based methods.
@@ 2.1 Mitigating Catastrophic Forgeting
performs continued training in a shared vector space, controlling learning process by adding regularization term on loss or applying specific optimization designs. Notabley, O-LoRA Wang et al. (2023b) introduce regularization with previous continual learned parameters for reducing interference in the multi-stage training setting. Our proposed CLoRA imposes orthogonal regularization similar to O-LoRA, but the regularization matrix is not restricted to be the previous learned parameter, thus CLoRA can be used for one-stage continued training whereas O-LoRA not.
@@ 3.2 Problem Definition
Catastrophic forgetting menifest as performance decline on tasks from previous domain when training on new domain. In this work, we aim to mitigate catastrophic forgetting in LLM finetuning and continual learning settings.
3.2.1 LLM Finetuning
In this setting, we conduct experiments on one-stage LLM finetuning, To evaluate this, we conduct both in-domain tasks (demonstrating the effectiveness of training) and out-domain tasks (from previous domain, indicating the degree of forgetting) for LLM finetuning. Specifically, we finetune a base LLM on one training dataset, then take in-domain and out-domain evaluations. Note that there is no clear domain specific for base LLMs, but benchmarks exist for evaluating the ability of LLMs on wide range of domains Gao et al. (2024), and we take those with minimal overlap with training data for out-domain evaluation.
3.2.2 Continual Learning
Continual learning focuses on developing learning algorithms to accumulate knowledge on non-stationary dataWang et al. (2023b). In this setting, we conduct experiments for multi-stage finetuning. Specifically, we finetune the model on a sequence of tasks $D_{1},\dots,D_{t}$, where each task $D_{t}$ contains a pair of train and test datasets $D_{t}=(D^{train}_{t},D^{test}_{t})$. The $t$-th model with finetuned sequentially on $D^{train}_{1},\dots,D^{train}_{t-1}$ is tested over all previous test datasets $D^{test}_{1},\dots,D^{test}_{t-1}$.
@@ 4 Method
For the CLoRA regularization matrices, following the priciple of Occam’s Razor, we adopt the most simple random initialization here. For uniform regularization over each row in regularization matrices, we suggest using orthogonal initialization. Specifically, for regularization matrix $P\in\mathbb{R}^{m\times k}$, $||P[:,i]||=1$ for every $i$, and $P[:,i]P[:,j]=0$ for $i\neq j$.
5 Experiments and Analysis
5.1 One-Stage LLM Finetuning
In this section, we conduct experiments on one-stage LLM finetuning to evaluate our proposed CLoRA as a parameter-efficient finetuning method. We aim to answer the following research questions,
  • RQ1: Does CLoRA performs effectively as a parameter-efficient finetuning method for LLMs with catastrophic forgetting mitigating?
  • RQ2: How the size of regularization matrix $k$ influence the performance of CLoRA? Does it differs across tasks?
@@ 5.1 One-Stage LLM Finetuning
5.1.1 Datasets and Tasks
Following previous works on PEFTLiu et al. (2024); Wang et al. (2024a), we conduct experiments on commonsense reasoning tasks and math tasks.
Commonsense Reasoning Setting
We use Commonsense170K Hu et al. (2023) for finetuning. For in-domain evaluation, eight commonsense reasoning datasets are used, including BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC-e, ARC-c Clark et al. (2018), and OBQA Mihaylov et al. (2018). The tasks are formulated as multiple-choice problem, and we report accuracy based on the last checkpoint.
For out-domain evaluations, BIG-Bench-Hard Suzgun et al. (2022) and MMLU-Pro Wang et al. (2024b) are used. These benchmarks encompass challenging subsets of tasks across a wide range of domains and are widely employed for evaluating the capabilities of LLMs. Additionally, they include samples that are more complex than those in our training data, ensuring minimal overlap. We use lm-eval Gao et al. (2024), available with MIT License, for reporting out-domain evaluation.
Math Setting
We use MetaMathQA Yu et al. (2024) for finetuning, which contains 395K samples augmented from the training set of GSM8K Cobbe et al. (2021) and MATHHendrycks et al. (2021). We use test set of GSM8K and MATH for evaluation and report the results on the last checkpoint.
5.1.2 Comparison Methods
@@
  • CLoRA: Our proposed CLoRA method, with random initialized regularization matrix.
In-domain Out-domain
Method BQ PQ SQ HS WG ACe ACc OQ Avg. BBH MMLU
LLaMA2-7b - - - - - - - - - 34.91 18.56
LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9 26.69 14.46
DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5 28.24 11.67
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8 29.54 11.33
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0 25.14 17.74
LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8 26.90 14.58
LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8 26.73 11.54
LoRA-L2 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8 32.93 16.59
CLoRA-k128 72.7 84.1 77.7 91.6 83.0 85.3 69.9 81.6 80.7 30.82 12.07
CLoRA-k256 71.3 83.2 79.1 92.4 83.2 84.5 71.0 81.0 80.7 31.92 17.81
CLoRA-k512 72.8 83.0 79.5 93.0 83.9 85.7 73.0 84.8 82.0 34.32 17.00
CLoRA-k1024 73.3 84.8 79.6 91.1 86.1 86.9 73.1 85.6 82.6 36.49 19.52
CLoRA-k2048 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7 38.67 20.59
@@
Table 2: Results for our proposed CLoRA and baselines for in-domain commonsense reasoning evaluations and out-domain LLM benchmarks, with accuracy scores (%) reported. Bold font indicates the highest performance for each task across all compared PEFT methods.
@@
5.1.3 Experimental Configuration
We use the same base LLM choice LLaMA-2-7B Touvron et al. (2023) and hyperparameter configurations as Wang et al. (2024a). Details are listed in Appendix A. Notably, we use 32 (commonsense reasoning) and 64 (math) for updating matrix rank $r$ as default for all methods if not explicitly specified. For the size of CLoRA regularization matrix, we select $k$ in $[128,256,512,1024,2048]$ for commonsense reasoning and $[64,128,256]$ for more challenging math setting. For LoRA-L2, 1e-5 is used for weighting of L2 regularization, we note that 1e-4 is also tested, but too large for getting effective finetuning. We report results finetuned on LLaMA-2-7B here, more results are listed in Appendix A.
5.1.4 Main Results (RQ1)
For commonsense reasoning setting, we report the results of in-domain evaluation and out-domain LLM benchmarks in Table 2. The results demonstrate that CLoRA outperforms on all datasets, surpassing the best baseline for in-domain evaluation by an average accuracy of 2.9 points. Results for math setting also demonstrate the superority of CLoRA over previous LoRA baselines (Table 3).
These outcomes suggest that, although primarily proposed for mitigating catastrophic forgetting, CLoRA also serves as an effective PEFT method. We attribute this to the nature of LLM finetuning, which is an instance of transfer learning. The performance of LLM finetuning is strongly correlated with the base model’s ability, when catastrophic forgetting occurs during training, the base model’s strength may diminish. Therefore, we claim that a method with effective capacity-forgetting balancing would exhibit strong effectiveness in LLM finetuning.
For out-domain evaluation, results show that all baselines underperform the base model, highlighting the severe issue of catastrophic forgetting in this setup. Notably, our proposed CLoRA not only outperforms all baselines by a significant margin but also surpasses the base model’s performance. We attribute this to CLoRA’s effective capacity-forgetting balancing, which enables the extraction of generally useful knowledge from the commonsense reasoning training dataset.
The superior performance in both in-domain and out-domain evaluations demonstrates that CLoRA serves as an effective parameter-efficient finetuning method with catastrophic forgetting mitigation.
The superior performance in both in-domain and out-domain evaluations demonstrates that our proposed CLoRA serves effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating. Thus, we answer RQ1.
5.1.5 Evaluating for different CLoRA $k$ (RQ2)
The size of the regularization matrix $k$ is a crucial hyperparameter in CLoRA, balancing the trade-off between model capacity and the degree of forgetting. We focus here on how $k$ influence the performance of finetuning LLM with CLoRA, and investigate whether the optimal $k$ is consistent across tasks.
In commonsense reasoning setting, results show that larger $k$ leads to better performance in both in-domain and out-domain evaluations (Table 2). In math setting, unlike the upward trend in commonsense reasoning setting, performance decreases when $k$ exceeds 128(Table 3). We attribute this discrepancy to the complexity of math tasks, which require greater model capacity during finetuning.
Emperical results support the intuitive claim that larger $k$ imposes more restrictions on updates, which helps mitigating catastrophic forgetting but potentially limiting finetuning model capacity and harming performance.
Thus, we answer RQ2 by demonstrating that the optimal $k$ depends on task complexity. Notably, our proposed CLoRA provides flexibility in balancing capacity and forgetting by adjusting $k$, we suggest choosing a smaller $k$ for more challenging tasks.
5.1.6 Understanding Capacity-Forgetting Balancing(RQ3)